The quadratic equation x^2 - 8x = -20 is to be solved by completing the square.
Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

I'll help you set it up.
take half of the coefficient of x (-8)
8/2=4
square it
16 add to both sides
x^2-8x+16=-4

Answer by checkley77(12844)

x^2 - 8x = -20
x^2-8x+16=-20+16
(x-4)^2=-4
x-4=+-(sqrt-4)
x-4=+-2i
x=4+-2i ans.